ganeshie8 (ganeshie8):

Factor \((111111)_2\) in binary

ganeshie8 (ganeshie8):

\[(111111)_2 = (63)_{10}\]

OpenStudy (kainui):

Ooooh the sum of digits of the factors is one more than the sum of digits of the original number.

ganeshie8 (ganeshie8):

This is related to a proof of mersenne numbers where the simplicity of factorization in binary is exploited for proving a mersenne number is not a mersenne prime im still trying to understand how the factorization in binary is easy let me tag the video..

OpenStudy (anonymous):

111111=1+10+100+1000+10000+100000 so seems prime to me >.<

ganeshie8 (ganeshie8):

https://www.youtube.com/watch?v=NrL-9XPqXq8#t=976

OpenStudy (anonymous):

haha i think its abut having this form (1111....11)(1000.....0001)

ganeshie8 (ganeshie8):

for some reason the alternative proof using below factorization looks lot simpler to me \[2^{pq}-1 = (2^p-1)((2^{p})^{q-1} +(2^{p})^{q-2}+ \cdots + 1 )\] how is factorization in binary better than this ? i feel im not seeing something obvious here..

OpenStudy (kainui):

Maybe it lends itself to some kind of byte code algorithm that happens real fast?

OpenStudy (anonymous):

lets consider 2^p-1 it would be of the form 1111...1 up to n 1's nw assume it have divisor of the form (1000000...01) samw number of digits as the origion and the other is (1111..1) 1 digit less

OpenStudy (kainui):

I'm still watching the video and he just finished talking bout odd semiperfect numbers.

ganeshie8 (ganeshie8):

Kai skip to 15th minute

ganeshie8 (ganeshie8):

@Marki im following u

OpenStudy (anonymous):

the thing is 0...0000001111111.....1 x 1000000........00000001 --------------- 0...00000111111.........1 + 1...11111000000.........0 ---------------------- 1...11111111111..........1

OpenStudy (kainui):

Ohhhh this is like that problem with elevens that we were looking at several months back.

ganeshie8 (ganeshie8):

1001 111 ---------- 1001 1001 1001 ------------- 111111 something like this ? but how do u know that the three 1's can be pulled out ?

ganeshie8 (ganeshie8):

111111 = 111(1001)

OpenStudy (kainui):

You know you can pull them out if there's a composite number of them I think, since it gives you a pattern that repeats

OpenStudy (anonymous):

so for 111111 it has 6 digits lets assume first divisor have half numbers of 1's digits 111 and the other of the form 1001 , so it could work lets see other example , shall we ?

ganeshie8 (ganeshie8):

lets try 2^8-1

ganeshie8 (ganeshie8):

yeah there is some pattern by which it can be factored i guess but i also think we need to know the factorization of 8 = 2*2*2 upfront hmm

ganeshie8 (ganeshie8):

\[2^8-1 = (11111111)_2\]

OpenStudy (kainui):

It might be nice to see how a composite number shows you how to factor it \[\large 2^{12}-1 = (111111111111)_2\] twelve factors into 1 (shown), 2, 3, 4, and 6 so we can factor it to look like these: \[11*10101010101 \\ 111*1001001001 \\ 1111*100010001 \\ 111111*1000001 \] See 2,3,4, and 6?

OpenStudy (kainui):

Here's all factorizations (first and last shown for completeness) 1, 2, 4, 8: \[2^8-1 \\ 1*11111111 \\ 11*1010101 \\ 1111 * 10101 \\ 11111111*1\]

OpenStudy (anonymous):

well 2^8 and 2^12 are already composite lets try when 2^p as p prime

ganeshie8 (ganeshie8):

I think it makes sense now xD \[2^8-1 \\ 1*11111111 \\ 11*1010101 \\ 1111 * 10\color{red}{0}01 \\ 11111111*1\]

OpenStudy (kainui):

Whoops, well you got the point lol. Just make a space that's n-1 of 0's between the 1's so that they fall into place. When 2^p is prime, then you can't factor it nicely I think, so it helps narrow it down. Interesting! I see now!

ganeshie8 (ganeshie8):

Got it completely ! thanks marki and kai :)

OpenStudy (anonymous):

wellif n=p1...pk then 2^n-1 divides 2^pk-1 :P

OpenStudy (anonymous):

thats why all tries works >.<

OpenStudy (anonymous):

lets try some prime number >.< like 2^19-1

OpenStudy (anonymous):

111....1 (19 digits ) so we wont be able to divide it which means it might be prime :)

OpenStudy (anonymous):

hmm nvm

OpenStudy (kainui):

I see so the only primes of the form 2^n-1 have to have n be prime. Strange

OpenStudy (kainui):

It's kind of like a geometric argument lol

OpenStudy (anonymous):

yes kai :) u can prove it anyway lol i guess ganesh and i proved it before hmm

OpenStudy (kainui):

I wonder if we can do a similar sort of argument in different bases. For instance, n!-1 in a factorial base?

OpenStudy (kainui):

In a factorial base, each digit holds a different amount, so 1=1!, 10=2!, 100=3!, 1000=4!, etc... so correspondingly, \[\Large 4! -1 = 321_{factorial} = 3*3!+2*2!+1*1!=23\]

OpenStudy (anonymous):

interesting >.< save this question lol

OpenStudy (anonymous):

oh wait

OpenStudy (kainui):

Interesting, here's an identity I just found \[\Large n!-1 = \sum_{k=1}^{n-1}k*k!\]

OpenStudy (anonymous):

u mean E base?

OpenStudy (kainui):

Yeah I think, where e is represented as a repeating decimal?

OpenStudy (anonymous):

not sure first u mean by ! repeated 10 but then u define 4! as 24 >.<

OpenStudy (kainui):

Oh I see how that's confusing, so I should have written: \[1_f = 1! \\ 10_f = 2! \\ 100_f = 3! \\ 1000_f=4!\] So each digit place has a different number of digits it can contain. So I'll show you how to count to 10 in this base: \[1,10,11,20,21,100,101,110,111,120\]

OpenStudy (anonymous):

can u define it in please >.< like write full definition

OpenStudy (kainui):

I don't know how uhhhh I don't know what it's called I just found it on my own.

OpenStudy (kainui):

\[1_f=1_{10} \\ 10_f=2_{10} \\ 11_f=3_{10} \\ 20_f=4_{10} \\ 21_f=5_{10} \\ 100_f=6_{10} \\ \]

OpenStudy (kainui):

So the rightmost digit only holds 0 and 1, the next digit holds 0,1, and 2, the next digit holds 0, 1, 2, and 3, etc... So each 1000_f is a factorial, so if you see there 6=100f and 2=10f because 2 is 2! and 6 is 3! if that makes sense.

OpenStudy (anonymous):

so 5!=100000f 5!-1=4321f

OpenStudy (anonymous):

1 10 11 20 21 100 101 110 111 120 121 200 201 210 211 220 221 300 301 310 311 320 321 1000 :D

OpenStudy (anonymous):

ok got it

OpenStudy (kainui):

Yeah! I just played around with Java and found that 12!-1 is a prime number 479001599. so it looks like 11.10.987654321 (weird since we have places that hold more than 10 in it haha XD

OpenStudy (anonymous):

i see 12!=1000 000 000 000 12!-1=111097654321 hmm

OpenStudy (anonymous):

so if 12!-1 is prime in f then its prime in real :O

OpenStudy (anonymous):

that reduce lots of work >.<

OpenStudy (anonymous):

so nw u can check numbers like 167574365! -1 :D

OpenStudy (anonymous):

whats the largest number u can check ?

OpenStudy (kainui):

Some primes, but I don't know how to efficiently check them, it's not like binary so I'm trying to find a different trick that will work. But here are ones we know, maybe we can find a pattern.\[21_f = 5 \\ 321_f = 23 \\ 54321_f = 719 \\ 13,12,11,10,987654321_f= 87178291199\] So specifically when n=3,4,6,12,14 are known primes so far. Plus we have that series formula above that might help us.

OpenStudy (anonymous):

ur series formula hmm i still feel dull about it u mean (n!-1) base 10= (...) in base f ?

OpenStudy (kainui):

Oh that's in base 10. Sorry for the confusion XD But the idea came from looking at just the digits: \[\Large (4!-1)_{10} = 321_f = (3*3!+2*2!+1*1!)_{10}\]

OpenStudy (anonymous):

so for n=6 (6!-1) base 10 = 1+20+300+4000+50000

OpenStudy (kainui):

Right, in base f.

OpenStudy (anonymous):

yes yes and to convert to 10 again 1*1!+2*2!+....+5*5! logic of bases :P

OpenStudy (anonymous):

brilliant , what else ?

OpenStudy (kainui):

Yeah haha you understand now, awesome. XD Ok so I should show you how I first came to this before, it was in seeing: \[\Large e = \frac{1}{0!}+\frac{1}{1!}+\frac{1}{2!}+\frac{1}{3!}+... \\ \Large e = 10. \bar 1_f\]

OpenStudy (anonymous):

i see !

OpenStudy (anonymous):

nw this shows e is irrational in neat way :O

OpenStudy (anonymous):

kai , write a paper about this seriously !

OpenStudy (kainui):

Haha I think someone else has probably found this before me. Ok so back to the problem though, I think if there is an even number of digits like \[\Large 4321_f = (4*4!+3*3!+2*2!+1*1!)_{10} =\\ \Large [(4^2+3)3! + (2^2+1)1!]_{10} \] maybe we can see something in the series?

OpenStudy (anonymous):

never seen it with f base ?

OpenStudy (anonymous):

n!-1= sum ((k^2+(k-1))(k-1)! ) from k=1 to n-1 ?

OpenStudy (kainui):

\[\Large n!-1 = \sum_{k=1}^{n-1} [(2k)^2+(2k-1)](2k)! \\ (for \ n \ odd)\]

OpenStudy (anonymous):

note that ohkk imissed it haha

OpenStudy (kainui):

wait almost

OpenStudy (kainui):

\[\Large n!-1 = \sum_{k=1}^{n-1} [(2k)^2+(2k-1)](2k-1)! \\ (for \ n \ odd)\]

OpenStudy (kainui):

no wait one more thing... XD

OpenStudy (anonymous):

:P

OpenStudy (kainui):

\[\Large n!-1 = \sum_{k=1}^{(n-1)/2} [(2k)^2+(2k-1)](2k-1)! \\ (for \ n \ odd)\]

OpenStudy (kainui):

And I guess there's also another formula for even n as well where you just pull the 1 out. #_# but I'm sleepy not used to being up this late haha.

OpenStudy (anonymous):

haha ok, gn anyway ! that was fun

OpenStudy (kainui):

Maybe we can try to figure out some thing about using them to find prime numbers or something in this funny base haha. Well, goodnight XD

OpenStudy (anonymous):

lets check this one (10000! -1) , in ur method it have 10000 numbers of digits but in real 35660